# Structural modulation, physical properties, and electronic band structure of the kagome metal UCr6Ge6

### v1

@@ III.3 Heat capacity and band structure
The heat capacity of UCr6Ge6 plotted as (Fig. 5), is mostly featureless but exhibits two bumps near 30 K attributed to an impurity phase (Fig. S3 [supplement]). A fit of between K yields mJ mol-1 K-2 and mJ mol-1 K-4. The large Sommerfeld coefficient of UCr6Ge6 is comparable to that reported for LuFe6Ge6, YbFe6Ge6, and Lu1-xFe6Sn6 ( mJ mol-1 K-2 [avila2005direct, lyu2024anomalous, shi2025weak]), though smaller than that of YbV6Sn6 ( mJ mol-1 K-2 [guo2023triangular]) and YbMn6Sn6 ( mJ mol-1 K-2 [lou2025strongly]). Additionally, the Debye temperature of UCr6Ge6 is determined to be =444(4) K, where is the number of atoms per formula unit (13) and is the gas constant. is consistent with the larger values (calculated with DFT) for Cr6Ge6 compounds with lighter elements [romaka2024structure]. As a comparison, the heat capacity of LuCr6Ge6 was also collected (Fig. S4(a) [supplement]). The fit Sommerfeld coefficient and Debye temperature were mJ mol-1 K-2 and K for LuCr6Ge6. The Sommerfeld coefficient of LuCr6Ge6 is similar to that of YCr6Ge6 (66 mJ mol-1 K-2), which has electronic bands near associated with the chromium kagome lattice [ishii2013ycr6ge6]. The similarity suggests that the larger electronic heat capacity of UCr6Ge6 may be driven by band structure features associated with not only chromium but also additional uranium 5 states that increase the DOS near .
@@
To assess whether the electronic heat capacity is influenced by chromium kagome flatbands and uranium 5 states, the electronic band structure was calculated for an ordered UCr6Ge6 cell with uranium 5 electrons in the valence [Fig. 6(a)] and for LuCr6Ge6 [Fig. 6(b)]. Additional band structures for each material with an effective Hubbard are provided in the Supplemental Material [supplement]. The calculated bands are shown with colors indicating weight from uranium/lutetium (teal), chromium (red), or germanium (pink) states. The general features of the UCr6Ge6 band structure are, unsurprisingly, similar to those reported for other 166s [thomas2025uv6sn6, riedel2025magnetic, hu2022tunable, ortiz2024stability, kim2023infrared], but for UCr6Ge6, flatbands typically associated with the kagome lattice along the – path sit below . The relative shift also places a flat valence band with hybrid uranium and chromium character along the – Brillouin zone path close to , further demonstrating the band tuning possibilities provided by expanding the 166 family to 5 systems.
@@
Comparing the UCr6Ge6, LuCr6Ge6, and previously published YCr6Ge6 [ishii2013ycr6ge6] band structures reveals that the Cr6Ge6 structures are conducive to a significant DOS at . The total DOS at for a UCr6Ge6 formula unit is 21.64 states/eV (13.79 states/eV for LuCr6Ge6). The equation
@@
relates the Sommerfeld coefficient to the DOS at [] and the Boltzmann constant () for noninteracting electrons. The predicted Sommerfeld coefficient is 51.0 mJ mol-1 K-2 for UCr6Ge6 and 32.5 mJ mol-1 K-2 for LuCr6Ge6. The ratio of the experimental to predicted value is, therefore, 1.7 for UCr6Ge6 and 1.8 for LuCr6Ge6, indicating moderate enhancement. While the enhancement may indicate mass renormalization due to electron correlations, the increase may also stem from electron-phonon coupling. When an effective Hubbard is included in the band structure calculation to localize the uranium 5 electrons, the chromium DOS near is mostly unchanged, and the uranium weight at decreases to nearly zero (Fig. S5(a) [supplement]), resulting in a predicted Sommerfeld coefficient of 26.4 mJ mol-1 K-2.
III.4 Resistivity
@@
Figure 7(a) shows the zero-field longitudinal resistivity of UCr6Ge6 with current applied in the kagome plane () and perpendicular to it (). The residual resistivity ratio [(300 K)/(1.8 K)] is 1.6 for and 3.7 for . Both are lower than the reported ratios for other Cr6Ge6 crystals, the structures of which are less disordered [konyk2020electrical, romaka2024structure], but the values are similar to those of UNb6Sn6 (3.3 [riedel2025magnetic]) and UV6Sn6 (2-3 [patino2025incom, thomas2025uv6sn6]). Neither orientation’s resistivity contains a feature near 30 K, and the first derivative of both orientations only shows a broad hump at the inflection point of the curve (Fig. S6 [supplement]). The resistivity with current along the axis () is lower than the resistivity with current within the plane (), as is also observed for YCr6Ge6 [ishii2013ycr6ge6] and for several CoSn-structure compounds [meier2020flat]. Figure 7(b) shows the magnetoresistance (MR) up to 9 T for both current directions with a perpendicular applied magnetic field. The magnetoresistance was calculated using
@@
As with the zero-field, temperature-dependent resistivity, no transitions appear. The changes are small and monotonic for both orientations and are consistent with preferred conduction perpendicular to the kagome lattice.
@@ IV Conclusions
UCr6Ge6 crystallizes in a structure with an average Y0.5Co3Ge3-type unit cell and a modulation that lowers the space-group symmetry to monoclinic. The modulation [] of UCr6Ge6 is distinct from other 166 materials in that it contains an component and no component. The isotropic temperature and magnetic field dependence of the magnetization contrast with other Cr6Ge6 and uranium-containing 166 materials with localized magnetic moments. Instead, the shape of the high-temperature susceptibility and the small magnetization strongly suggest that UCr6Ge6 is Pauli paramagnetic, though a ground state singlet due to a CEF splitting of a U 5 multiplet cannot be ruled out. A transition at 30 K in the magnetic susceptibility and heat capacity are attributed to an impurity phase. The electronic heat capacity, quantified by a Sommerfeld coefficient of 86.5 mJ mol-1 K-2, is moderately enhanced in UCr6Ge6 possibly due to flatbands near the Fermi level or electron-phonon coupling. Electronic band structure calculations suggest that the kagome flatbands and uranium states may increase the DOS near . The anisotropic resistivity and magnetoresistance reveal a preference for conduction perpendicular to the kagome lattice, consistent with other CoSn-type and stuffed-CoSn-type materials. The moderately enhanced Sommerfeld coefficient, isotropic magnetic properties, and unique structural modulation distinguish UCr6Ge6 from other uranium-containing 166 compounds and demonstrate the promise of tuning electronic band structure features and magnetic ground states in 5 166 materials.

### v2

@@ III.3 Heat capacity and band structure
The heat capacity of UCr6Ge6 plotted as (Fig. 5), is mostly featureless but exhibits two bumps near 30 K attributed to an impurity phase (Fig. S3 [1]). A fit of between K yields mJ mol-1 K-2 and mJ mol-1 K-4. The large Sommerfeld coefficient of UCr6Ge6 is comparable to that reported for LuFe6Ge6, YbFe6Ge6, and Lu1-xFe6Sn6 ( mJ mol-1 K-2 [4, 35, 58]), though smaller than that of YbV6Sn6 ( mJ mol-1 K-2 [16]) and YbMn6Sn6 ( mJ mol-1 K-2 [33]). Additionally, the Debye temperature of UCr6Ge6 is determined to be =444(4) K, where is the number of atoms per formula unit (13) and is the gas constant. is consistent with the larger values (calculated with DFT) for Cr6Ge6 compounds with lighter elements [51]. As a comparison, the heat capacity of LuCr6Ge6 was also collected (Fig. S4(a) [1]). The fit Sommerfeld coefficient and Debye temperature were mJ mol-1 K-2 and K for LuCr6Ge6. The Sommerfeld coefficient of LuCr6Ge6 is similar to that of YCr6Ge6 (66 mJ mol-1 K-2), which has electronic bands near associated with the chromium kagome lattice [20]. The similarity suggests that the larger electronic heat capacity of UCr6Ge6 may be driven by band structure features associated with not only chromium but also additional uranium 5 states that increase the DOS near .
@@
To assess whether the electronic heat capacity is influenced by chromium kagome flatbands and uranium 5 states, the electronic band structure was calculated for an ordered UCr6Ge6 cell with uranium 5 electrons in the valence [Fig. 6(a)] and for LuCr6Ge6 [Fig. 6(b)]. Additional band structures for each material with an effective Hubbard are provided in the Supplemental Material [1]. The calculated bands are shown with colors indicating weight from uranium/lutetium (teal), chromium (red), or germanium (pink) states. The general features of the UCr6Ge6 band structure are, unsurprisingly, similar to those reported for other 166s [59, 50, 19, 44, 23], but for UCr6Ge6, flatbands typically associated with the kagome lattice along the – path sit below . The relative shift also places a flat valence band with hybrid uranium and chromium character along the – Brillouin zone path close to , further demonstrating the band tuning possibilities provided by expanding the 166 family to 5 systems.
@@
Comparing the UCr6Ge6, LuCr6Ge6, and previously published YCr6Ge6 [20] band structures reveals that the Cr6Ge6 structures are conducive to a significant DOS at . The total DOS at for a UCr6Ge6 formula unit is 21.64 states/eV (13.79 states/eV for LuCr6Ge6). The equation
@@
relates the Sommerfeld coefficient to the DOS at [] and the Boltzmann constant (). The predicted Sommerfeld coefficient is 51.0 mJ mol-1 K-2 for UCr6Ge6 and 32.5 mJ mol-1 K-2 for LuCr6Ge6. The ratio of the experimental to predicted value is, therefore, 1.7 for UCr6Ge6 and 1.8 for LuCr6Ge6, indicating moderate enhancement. When an effective Hubbard is included in the band structure calculation to localize the uranium 5 electrons, the chromium DOS near is mostly unchanged, and the uranium weight at decreases to nearly zero (Fig. S5(a) [1]), resulting in a predicted Sommerfeld coefficient of 26.4 mJ mol-1 K-2.
III.4 ARPES
High-resolution ARPES was used to probe the electronic structure of UCr6Ge6 directly to determine if the enhanced Sommerfeld coefficient observed is related to the presence of flat bands in the vicinity of the Fermi level. Constant energy contours in the - plane were measured and are presented in Fig. 7. As shown in Fig. 7(a), there is little structure at the Fermi surface within the first Brillouin zone, while spectral weight with little momentum variation is observed in neighboring zones. This is in sharp contrast to the Fermi surface observed in UV6Sn6, in which the prototypical six-fold symmetric kagome structure is clearly resolved [59]. However, a six-fold symmetric structure as well as triangular pockets about the point are observed in a constant energy contour recorded at eV, as displayed in Fig. 7(b). This suggests an electron band filling which shifts the Fermi level of order 300 meV in UCr6Ge6 relative to UV6Sn6.
To search for evidence of 3 Cr kagome flatbands near the Fermi level, valence band spectra were collected with photon energies of 92 eV (uranium Cooper minimum) and 108 eV (uranium 5-5, or O-edge, resonance condition), utilizing linear horizontal (LH) and linear vertical (LV) polarized light to account for matrix element effects. These energies correspond to values that lie close to the and the planes, respectively (Fig. S6 [1]). When probed with LH polarization, a clear flatband is observed near the Fermi level [Fig. 8(a)], the spectral weight of which is suppressed in the first Brillouin zone due to matrix element effects. With LV-polarized light, the band is also weakly observed in the first Brillouin zone [Fig. 8(c)]. As this measurement is performed at the Cooper minimum of uranium, we ascribe Cr character, likely from even-parity Cr orbitals, to these flat bands. The presence of a kagome flat band in the vicinity of the Fermi level is a strong candidate for the origin of the enhanced Sommerfeld coefficient observed in UCr6Ge6, contrasting with UV6Sn6, where a flat band near the Fermi level is absent.
Furthermore, in Fig. 8(b,d), we show valence band dispersions that enhance uranium character due to the uranium - resonance condition. In Fig. 8(b,d), there is enhanced momentum-independent spectral weight at the Fermi level characteristic of conduction–-electron () hybridization in heavy fermion systems. The larger spectral weight found in UCr6Ge6 compared to that of UV6Sn6 indicates that the electrons in UCr6Ge6 are more itinerant than in UV6Sn6. These results are consistent with the scenario of Pauli paramagnetism in UCr6Ge6 arising from itinerant electrons, while the more localized electrons in UV6Sn6 lead to magnetic order [59, 2].
III.5 Resistivity
@@
Figure 9(a) shows the zero-field longitudinal resistivity of UCr6Ge6 with current applied in the kagome plane () and perpendicular to it (). The residual resistivity ratio [(300 K)/(1.8 K)] is 1.6 for and 3.7 for . Both are lower than the reported ratios for other Cr6Ge6 crystals, the structures of which are less disordered [26, 51], but the values are similar to those of UNb6Sn6 (3.3 [50]) and UV6Sn6 (2-3 [2, 59]). Neither orientation’s resistivity contains a feature near 30 K, and the first derivative of both orientations only shows a broad hump at the inflection point of the curve (Fig. S7 [1]). The resistivity with current along the axis () is lower than the resistivity with current within the plane (), as is also observed for YCr6Ge6 [20] and for several CoSn-structure compounds [38]. Figure 9(b) shows the magnetoresistance (MR) up to 9 T for both current directions with a perpendicular applied magnetic field. The magnetoresistance was calculated using
@@
As with the zero-field, temperature-dependent resistivity, no transitions appear. The changes are small and monotonic for both orientations and are consistent with preferred conduction perpendicular to the kagome lattice. Additional magnetoresistance (, ) and Hall resistivity () data at 2, 10, 20, 30, and 50 K are presented in the Supplemental Material [1]. At higher temperatures, the magnetoresistance is approximately flat and near 0%, and the Hall resistivity indicates predominant electron carriers in a one-band approximation. A linear fit of to at 2 K results in a carrier concentration cm-3 (Fig. S9), where is the Hall coefficient and is the electron charge.
@@ IV Conclusions
UCr6Ge6 crystallizes in a structure with an average Y0.5Co3Ge3-type unit cell and a modulation that lowers the space-group symmetry to monoclinic. The modulation [] of UCr6Ge6 is distinct from other 166 materials in that it contains an component and no component. The isotropic temperature and magnetic field dependence of the magnetization contrast with other Cr6Ge6 and uranium-containing 166 materials with localized magnetic moments. Instead, the shape of the high-temperature susceptibility and the small magnetization strongly suggest that UCr6Ge6 is Pauli paramagnetic, though a ground state singlet due to a CEF splitting of a U 5 multiplet cannot be ruled out. A transition at 30 K in the magnetic susceptibility and heat capacity are attributed to an impurity phase.
The electronic heat capacity, quantified by a Sommerfeld coefficient of 86.5 mJ mol-1 K-2, is moderately enhanced in UCr6Ge6, possibly due to flatbands near the Fermi level. Electronic band structure calculations suggest that the kagome flatbands and uranium states may increase the DOS near .
The ARPES study reveals that the Fermi level is shifted 300 meV in UCr6Ge6 relative to UV6Sn6. This shift results in Cr 3 flatbands now residing near the Fermi level, giving a large density of states from which the enhanced Sommerfeld coefficient can be explained. Furthermore, the U states exhibit hybridization, indicating more itinerant character in UCr6Ge6, consistent with a Pauli paramagnetic state.
The anisotropic resistivity and magnetoresistance reveal a preference for conduction perpendicular to the kagome lattice, consistent with other CoSn-type and stuffed-CoSn-type materials. The moderately enhanced Sommerfeld coefficient, isotropic magnetic properties, and unique structural modulation distinguish UCr6Ge6 from other uranium-containing 166 compounds and demonstrate the promise of tuning electronic band structure features and magnetic ground states in 5 166 materials.
Acknowledgements.
Crystal growth, physical property measurements, and magnetic property measurements were supported by the U.S. Department of Energy, Office of Basic Energy Sciences, Division of Materials Science and Engineering project “Quantum fluctuations in narrow band systems.” Structural characterization was also supported by the Laboratory Directed Research and Development program. C.S.K. and P.A.E.M. gratefully acknowledges the support through the LANL/LDRD Program and the G. T. Seaborg Institute. We gratefully acknowledge Diamond Light Source for access to Beamline I05 (Proposal SI41855-1), which contributed to the results presented here. Scanning electron microscope and energy dispersive X-ray measurements were performed at the Electron Microscopy Lab and supported by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy Office of Science. The electronic structure calculations were supported in part by the Center for Integrated Nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resources. Additional computations were performed at the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award ERCAP0028014.References
- [1] Note: See Supplemental Material at [URL will be inserted by publisher] for additional information, including Refs. [36, 21]. Cited by: §II.1, §III.1, §III.1, §III.1, §III.2, §III.2, §III.3, §III.3, §III.3, §III.4, §III.5, §III.5.
- [2] (2025) Incommensurate and commensurate antiferromagnetic orders in the kagome compound UV6Sn6. Physical Review B 111 (17), pp. 174432. Cited by: §I, §III.1, §III.2, §III.4, §III.5.
- [3] (2022) Charge density wave in kagome lattice intermetallic ScV6Sn6. Physical Review Letters 129 (21), pp. 216402. Cited by: §I.
- [4] (2005) Direct observation of Fe spin reorientation in single-crystalline YbFe6Ge6. Journal of Physics: Condensed Matter 17 (43), pp. 6969. Cited by: §I, §III.3.
- [5] (1994) Magnetic properties of RCr6Ge6 compounds. Journal of Alloys and Compounds 205 (1-2), pp. 77–80. Cited by: §III.2.
- [6] (2025) SAINT V8.42. Cited by: §II.1.
- [7] (1981) Intermetallische phasen mit B35-Überstruktur und verwandtschaftsbeziehung zu LiFe6Ge6. Zeitschrift für anorganische und allgemeine Chemie 482 (11), pp. 40–48. Cited by: §I, §III.1.
- [8] (1999) Magnetic properties of RMn6Sn6 (R=Tb, Ho, Er, Tm, Lu) single crystals. Journal of Magnetism and Magnetic Materials 207 (1-3), pp. 78–94. Cited by: §I.
- [9] (2021) Anisotropically large anomalous and topological Hall effect in a kagome magnet. Physical Review B 104 (16), pp. L161115. Cited by: §I.
- [10] (1998) Electron-energy-loss spectra and the structural stability of nickel oxide: An LSDA+U study. Physical Review B 57 (3), pp. 1505. Cited by: §II.2.
- [11] (2008) Origins of superstructure ordering and incommensurability in stuffed CoSn-type phases. Journal of the American Chemical Society 130 (26), pp. 8195–8214. Cited by: §III.1, §III.1.
- [12] (2012) Complex itinerant ferromagnetism in noncentrosymmetric Cr11Ge19. Physical Review B 85 (22), pp. 224405. Cited by: §III.2.
- [13] (2020) Competing magnetic phases and fluctuation-driven scalar spin chirality in the kagome metal YMn6Sn6. Science Advances 6 (51), pp. eabe2680. Cited by: §I.
- [14] (1994) UFe6Ge6: A new ternary magnetic compound. Journal of Alloys and Compounds 204 (1-2), pp. 59–64. Cited by: §I.
- [15] (2016) The Cambridge structural database. Structural Science 72 (2), pp. 171–179. Cited by: §II.1.
- [16] (2023) Triangular Kondo lattice in YbV6Sn6 and its quantum critical behavior in a magnetic field. Physical Review B 107 (20), pp. 205151. Cited by: §I, §III.3.
- [17] (2017) A facility for the analysis of the electronic structures of solids and their surfaces by synchrotron radiation photoelectron spectroscopy. Review of Scientific Instruments 88 (1). Cited by: §II.1.
- [18] (2024) Phonon promoted charge density wave in topological kagome metal ScV6Sn6. Nature Communications 15 (1), pp. 1658. Cited by: §I.
- [19] (2022) Tunable topological Dirac surface states and van Hove singularities in kagome metal GdV6Sn6. Science Advances 8 (38), pp. eadd2024. Cited by: §III.3.
- [20] (2013) YCr6Ge6 as a candidate compound for a kagome metal. Journal of the Physical Society of Japan 82 (2), pp. 023705. Cited by: §I, §I, §III.3, §III.3, §III.5.
- [21] (2000) SUBGROUPGRAPH: a computer program for analysis of group–subgroup relations between space groups. Applied Crystallography 33 (4), pp. 1190–1191. Cited by: §III.1, 1.
- [22] (2025) Small Seebeck in A15-type Cr3Ge single crystals. Europhysics Letters. Cited by: §III.2.
- [23] (2023) Infrared probe of the charge density wave gap in ScV6Sn6. Physical Review B 108 (20), pp. 205118. Cited by: §I, §III.3.
- [24] (2006) High-field magnetization of RMn6Sn6 compounds with R=Gd, Tb, Dy and Ho. Journal of Alloys and Compounds 408, pp. 169–172. Cited by: §I.
- [25] (1980) ESCA and magnetic studies of the Cr-Ge system. Journal of Magnetism and Magnetic Materials 20 (1), pp. 99–106. Cited by: §III.2.
- [26] (2020) Electrical transport properties of RCr6Ge6 (R=Y, Gd, Tb, Dy, Lu) compounds. Visnyk of the Lviv University Chemistry Series, pp. 107–113. External Links: Document Cited by: §I, §III.1, §III.5.
- [27] (2015) Comparison of silver and molybdenum microfocus X-ray sources for single-crystal structure determination. Applied Crystallography 48 (1), pp. 3–10. Cited by: §II.1.
- [28] (1996) Efficient iterative schemes for total-energy calculations using a plane-wave basis set. Physical Review B 54 (16), pp. 11169. Cited by: §II.2.
- [29] (1993) Molecular dynamics for open-shell transition metals. Physical Review B 48 (17), pp. 13115. Cited by: §II.2.
- [30] (1999) From ultrasoft pseudopotentials to the projector augmented-wave method. Physical Review B 59 (3), pp. 1758. Cited by: §II.2.
- [31] (2025) High-field magnetic phase diagrams of the Mn6Sn6 (= Gd–Tm) kagome metals. Physical Review B 111 (5), pp. 054410. Cited by: §I.
- [32] (2025) Coexisting kagome and heavy fermion flat bands in YbCr6Ge6. External Links: 2509.04902 Cited by: §I, §I.
- [33] (2025) Strongly Entangled Kondo and Kagome Lattices and the Emergent Magnetic Ground State in Heavy-Fermion Kagome Metal YbV6Sn6. Physical Review Letters 135 (14), pp. 146902. Cited by: §I, §III.3.
- [34] (2026) Cooperative concurrence of and flat bands in kagome heavy-fermion metal YbCr6Ge6. External Links: 2601.05829 Cited by: §I.
- [35] (2024) Anomalous Hall effect and electronic correlation in a spin-reoriented kagome antiferromagnet LuFe6Sn6. Chinese Physics B 33 (10), pp. 107507. Cited by: §I, §III.3.
- [36] (1997) Magnetic properties of NdMn6Sn6 and SmMn6Sn6 compounds from susceptibility measurements and neutron diffraction study. Journal of Alloys and Compounds 252 (1-2), pp. 41–49. Cited by: §III.1, 1.
- [37] (2001) Macroscopic magnetic properties of the HfFe6Ge6-type RFe6X6 (X= Ge or Sn) compounds involving a non-magnetic R metal. Journal of Alloys and Compounds 325 (1-2), pp. 67–72. Cited by: §III.2.
- [38] (2020) Flat bands in the CoSn-type compounds. Physical Review B 102 (7), pp. 075148. Cited by: §III.5.
- [39] (2025) Pressure suppresses the density wave order in kagome metal LuNb6Sn6. Physical Review Materials 9 (8), pp. L082001. Cited by: §I.
- [40] (2023) Tiny Sc allows the chains to rattle: Impact of Lu and Y doping on the charge-density wave in ScV6Sn6. Journal of the American Chemical Society 145 (38), pp. 20943–20950. Cited by: §I.
- [41] (2019) Overview of the U3Ge5 family with = Ti, V, Cr, Mn, Zr, Nb, Mo, Hf, Ta and W: Nine new members, phase formation, stability, structural and physical properties and electronic structures. Journal of Solid State Chemistry 277, pp. 260–270. Cited by: §III.2, §III.2.
- [42] (1993) 155Gd Mössbauer effect and magnetic properties of GdMn6-xCrxGe6. Journal of Alloys and Compounds 198, pp. L1–L3. External Links: Document Cited by: §I.
- [43] (1992) Magnetic and electrical properties of U–Ge intermetallic compounds. Journal of the Physical Society of Japan 61 (1), pp. 293–299. Cited by: §III.2.
- [44] (2024) Stability frontiers in the AM6X6 kagome metals: the LnNb6Sn6 (Ln:Ce-Lu,Y) family and density-wave transition in LuNb6Sn6. Journal of the American Chemical Society. Cited by: §I, §III.3.
- [45] (1996) Generalized gradient approximation made simple. Physical Review Letters 77 (18), pp. 3865. Cited by: §II.2.
- [46] (2022) Highly anisotropic magnetism in the vanadium-based kagome metal TbV6Sn6. Physical Review Materials 6 (10), pp. 104202. Cited by: §I.
- [47] (2023) Frustrated charge order and cooperative distortions in ScV6Sn6. Physical Review Materials 7 (10), pp. 104201. Cited by: §I.
- [48] (2021) Electronic properties of the topological kagome metals YV6Sn6 and GdV6Sn6. Physical Review B 104 (23), pp. 235139. Cited by: §I.
- [49] (2024) New insight into tuning magnetic phases of Mn6Sn6 kagome metals. npj Quantum Materials 9 (1), pp. 42. Cited by: §I.
- [50] (2025) Magnetic order and physical properties of the kagome metal UNb6Sn6. Physical Review Materials 9, pp. 084401. Cited by: §I, §I, §III.1, §III.2, §III.3, §III.5.
- [51] (2024) Structure, bonding, and properties of Cr6Ge6 intermetallics (=Gd–Lu). Journal of Solid State Chemistry 338, pp. 124874. Cited by: §I, §III.1, §III.3, §III.5.
- [52] (1988) Electrical resistivity and magnetic susceptibility in the amorphous CrxGe1-x alloy system. Physical Review B 38 (16), pp. 11741. Cited by: §III.2.
- [53] (1983) Magnetic and electrical properties of CrGe and Cr11Ge8. Journal of the Physical Society of Japan 52 (5), pp. 1807–1813. Cited by: §III.2.
- [54] (1997) Atomic disorder and canted ferrimagnetism in the TbCr6Ge6 compound. a neutron study. Journal of Alloys and Compounds 255 (1-2), pp. 67–73. Cited by: §I, §III.2.
- [55] (1997) Ferrimagnetism and disorder in the RCr6Ge6 compounds (R=Dy, Ho, Er, Y): A neutron study. Journal of Alloys and Compounds 256 (1-2), pp. 92–96. Cited by: §I, §III.2.
- [56] (2015) Crystal structure refinement with SHELXL. Acta Crystallographica Section C 71 (1), pp. 3–8. Cited by: §II.1.
- [57] (2015) SHELXT–integrated space-group and crystal-structure determination. Acta Crystallographica Section A 71 (1), pp. 3–8. Cited by: §II.1.
- [58] (2025) Weak low-temperature ferromagnetism and linear magnetoresistance in Lu0.75Fe6Sn6 with a disordered HfFe6Ge6-type structure. Journal of Magnetism and Magnetic Materials 615, pp. 172793. Cited by: §I, §III.3.
- [59] (2025) Unusual magnetism in new kagome material UV6Sn6. npj quantum materials 10, pp. 66. Cited by: §I, §III.1, §III.1, §III.2, §III.3, §III.4, §III.4, §III.5.
- [60] (2002) Magnetotransport of compounds in the U–Ge system. Philosophical Magazine B 82 (7), pp. 805–824. Cited by: §III.2.
- [61] (2023) Dynamics and resilience of the unconventional charge density wave in ScV6Sn6 bilayer kagome metal. Communications Materials 4 (1), pp. 103. Cited by: §I.
- [62] (1985) The Cr–U (Chromium–Uranium) system. Bulletin of Alloy Phase Diagrams 6, pp. 425–429. Cited by: §III.2.
- [63] (2001) Crystallographic and magnetic properties of TbFe6Ge6-xGax compounds (0.53.5). Journal of alloys and compounds 329 (1-2), pp. 8–21. Cited by: §III.1.
- [64] (2006) Filling the CoSn host-cell: the HfFe6Ge6-type and the related structures. Zeitschrift für Kristallographie-Crystalline Materials 221 (5-7), pp. 511–520. Cited by: §III.1.
- [65] (2005) Crystal structure, 57Fe Mössbauer spectroscopy and magnetization of UxFe6Sn6 (00.6). Intermetallics 13 (5), pp. 490–496. Cited by: §I.
- [66] (2020) Experimental observation of electronic structures of kagome metal YCr6Ge6. Chinese Physics Letters 37 (8), pp. 087102. Cited by: §I.
- [67] (2024) Preparation, crystal structure, and properties of the kagome metal ThV6Sn6. Inorganic Chemistry 63 (49), pp. 23288–23295. Cited by: §I.
- [68] (2024) Crystal growth, magnetic and electrical transport properties of the kagome magnet Cr6Ge6 (= Gd–Tm). Chinese Physics B 33 (7), pp. 077501. Cited by: §I, §III.2.
- [69] (2024) Tuning charge density wave of kagome metal ScV6Sn6. New Journal of Physics 26 (5), pp. 052001. Cited by: §I.
- [70] (1968) Magnetic susceptibility and electrical conductivity of the highest chromium germanide. Soviet Physics Journal 11 (5), pp. 23–25. Cited by: §III.2.